The function x = (4.3 m) cos[(3πrad/s)t + π/5 rad] gives the simple harmonic motion of a body. At t = 2.2 s, what are the (a) displacement, (b) velocity, (c) acceleration, and (d) phase of the motion? Also, what are the (e) frequency and (f) period of the motion?
Answer #1

displacemnt,
x = 4.3* cos (3πt + π/5)

velocity = dx/dt
v = - 4.3* 3π sin (3πt + π/5)
   = -40.53 sin (3πt + π/5)

acceleration = dv/dt
a = -40.53*3π cos (3πt + π/5)
    = -382 cos (3πt + π/5)

a)
x = 4.3* cos (3πt + π/5)
put t= 2.2 s
x = 4.3* cos (3π*2.2 + π/5)
    =4.3 * cos (21.36 rad)
    = -3.48 m
b)
v = -40.53 sin (3πt + π/5)
put t= 2.2 s
v = -40.53 sin (3π*2.2 + π/5)
    = -40.53 sin (21.36 rad)
   = - 23.9 m/s
c)
a = -382 cos (3πt + π/5)
put t= 2.2 s
a = -382* cos (21.36 rad)
    = -308.4 m/s^2
d)
phase = 3πt + π/5
put t= 2.2 s
phase = 3π*2.2 + π/5
             = 21.36 rad
e)
w= 3π from given equation
use:
w= 2*π*f
2*π*f = 3*π
f = 3/2 Hz

f)
T = 1/f
    = 2/3 s